Please help I have solved part of this but I'm lost with the complete answer
Find the equation of the line Write in standard form with no fractions and a positive x term
Line through (-5,4) & (-6,2)
I get y=(2)x + (14) But how do I put that in standard form?
Thank you

Answer by venugopalramana(3286)

Please help I have solved part of this but I'm lost with the complete answer
Find the equation of the line Write in standard form with no fractions and a positive x term
Line through (-5,4) & (-6,2)
I get y=(2)x + (14) But how do I put that in standard form?
Thank you
Y-2 = [(2-4)/(-6+5)](X+6)=2X+12
Y=2X+14....
GOOD YOUR ANSWER IS CORRECT.
AS REGARDS STANDARD OR ,THERE ARE SEVERAL FORMS...THE MOST COMMON IS
1.SLOPE INTERCEPT FORM...Y = MX + C.......WHICH IS WHAT YOU HAVE GOT
Y=2X+14...SLOPE = M =2.........INTERCEPT=C= 14
2. LINEAR EQN. FORM.....AX+BY+C=0
2X-Y+14=0
I FEEL YOUR ANSWER IS GOOD TO BE LEFT ALONE